OpenStudy (anonymous):

What is the slope of the line that passes through (2, 3) and (4, 5)? b. What is the equation of the line with slope 4 and y-intercept at (0, 3)?

OpenStudy (anonymous):

wait soory i got the numbers wrong let me repost the question

OpenStudy (anonymous):

equation of line (y-y1)/(x-x1)= (y2-y1)/(x2-x1) where (y2-y1)/(x2-x1) is the slope of line 1. (x1=2, y1=3) and (x2=4, y2=5)? b. What is the equation of the line with slope 4 and y-intercept at (0, 3)? here y-3= 4(x-0)